The perimeter of a rectangular garden is 268 feet. If the length of the garden is 83 feet, what is its width?
Answer by checkley79(3341)

P=2L+2W
L=83
268=2*83+2W
268=166+2W
2W=268-166
2W=102
W=102/2
W=51 FT. ANS. FOR THE WIDTH.
PROOF:
268=2*83+2*51
268=166+102
268=268